OpenStudy (anonymous):

What is the value of x in the equation 4x+8y=40, when y=0.8 4.6 8.4 10 12

Mehek (mehek14):

put 0.8 in for y \(4x+8*0.8=40\\4x+6.4=40\\4x=33.6\) divide \(\dfrac{33.6}{4}\)

OpenStudy (anonymous):

Thanks